4. Find the square roots of the following numbers by the Prime Factorisation Method.  ii. 400  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


400 = 2×2×2×2×5×5×1
⇒ 400 = (2×2)×(2×2)×(5×5) ⇒400 = (2×2×5)×(2×2×5)⇒ 400 = (2×2×5)2⇒√400=2×2×5=20